# MoSECroT: Model Stitching with Static Word Embeddings for Crosslingual Zero-shot Transfer

### v1

@@ 3 MoSECroT Task Setting
The task setting is straightforward: given a PLM of a high-resource language (regarded as the source language) and static word embeddings of another language (possibly low-resource and regarded as the target language), the goal is to achieve zero-shot transfer by using the target language embeddings directly with the source language model via embedding layer stitching. This can be done by first applying an alignment between the source and target embedding spaces and subsequently swapping the embedding matrices of the PLM.
@@ 4 Methodology
Following Moschella et al. (2023), we build relative representations (RRs) for each token in the embedding space based on their similarities with anchor tokens in the respective language. Specifically, we compute the cosine similarity of the embedding of each token with the embedding of each anchor token. For example, in the source language, the similarity between token $x_{i}$ and anchor $a_{j}$ is calculated as follows:
@@
Note that the relative representation is sensitive to the order of the anchors, so the relative representation for each token is computed with the anchors in the same order. This computation results in a matrix $\boldsymbol{R}^{s}\in\mathbb{R}^{|V^{s}|\times|A|}$ for the source language and a matrix $\boldsymbol{R}^{t}\in\mathbb{R}^{|V^{t}|\times|A|}$ for the target language, where $|V^{s}|$ (resp. $|V^{t}|$) is the source-language (resp. target-language) vocabulary size and $|A|$ is the number of parallel anchors.
@@
The obtained relative representations are vectors in $\mathbb{R}^{|A|}$ for both source and target languages. This dimension does not suit the hidden dimension of the Transformer body of the source PLM. Therefore, we propose to map the relative representations of both source and target languages back to $\mathbb{R}^{D}$. Given $\boldsymbol{E}^{s}$ and $\boldsymbol{R}^{s}$ for source language (resp. $\boldsymbol{E}^{t}$ and $\boldsymbol{R}^{t}$ for target language), we compute the transformed embedding of any token $x_{i}$ from the source language (resp. any token $y_{i}$ from the target language) as follows:
@@
By far we project the target-language embeddings to $\mathbb{R}^{D}$ which suits the hidden dimension of the Transformer body of the source language. We can simply fine-tune the model ($\boldsymbol{F}^{s}$ and the Transformer body) on the source-language train set of a downstream task and then assemble a target-language model for zero-shot transfer, without training on the target language. To do this, we only need to swap the source-language embeddings $\boldsymbol{F}^{s}$ with target-language embeddings $\boldsymbol{F}^{t}$.
@@ 5.1 Setup
We evaluate the proposed method on two text classification datasets: Multilingual Amazon Reviews Corpus (Keung et al., 2020) and Taxi1500 (Ma et al., 2023). See Appendix C for details.
Apart from the standard weighting scheme illustrated in Section 4, we propose two more settings: one where we apply softmax over relative representation weights (in Embedding mapping step), and another using sparsemax (Martins and Astudillo, 2016). Compared to softmax, sparsemax produces sparse weight distributions, meaning more similarities are concentrated on fewer anchors. We conduct preliminary experiments to identify the optimal top-$k$ closest anchors $\in\{1,10,50,100\}$ and find that the results are best when using the top 50 anchors.
5.2 Baselines
We compare our method against three baselines:
Logistic Regression (LG)
mBERT
We fine-tune multilingual BERT (mBERT) (Devlin et al., 2019), using the English training data, and perform zero-shot predictions directly on the target language test data.
Least squares projection (LS)
We propose a straightforward approach to project target language embeddings into the same space as the English PLM embeddings. Specifically, we learn a transformation matrix $\boldsymbol{W}\in\mathbb{R}^{D^{t}\times D}$ by minimizing $||\boldsymbol{A}^{t}\boldsymbol{W}-\boldsymbol{A}^{s}||_{F}^{2}$, where $\boldsymbol{A}^{t}\in\mathbb{R}^{|A|\times D^{t}}$ is the embeddings of anchors in the target language and $\boldsymbol{A}^{s}\in\mathbb{R}^{|A|\times D}$ is the embeddings of anchors from the English PLM. We then project all target language embeddings using $\boldsymbol{W}$ and replace the BERT embedding layer with the resulting matrix.
@@ 5.3 Results
We present evaluation results of RRs with the proposed settings (5.1) and compare them with the baselines in Tables 1 and 2. Macro $F_{1}$ is used due to class imbalance in both datasets. We notice that the naive LS baseline is almost always beaten by the proposed method under multiple RR settings. The only exception is nds, where both LS and RRs perform badly. This observation is a strong indicator that RRs can better leverage the semantic similarity encoded in different types of embeddings than LS. Not very surprisingly, zero-shot with mBERT is effective for high-resource languages in both datasets but underperforms LS with large gaps on low-resource languages in Taxi1500, which indicates mBERT is not well-aligned across low-resource languages due to data sparsity. In contrast, for all five low-resource languages not seen by mBERT, RRs surpass mBERT, although the margin varies (ranging from +0.12 for sah to +0.01 for nds). On high-resource languages, none of the RR settings outperforms mBERT.
@@ 6 Analysis
We try to decrease the influence of potentially noisy anchor pairs by reducing the number of anchors to 3000 and 500777Original set contains about 6000 anchors. through random sampling, following the observation by Moschella et al. (2023) that uniform selection from an anchor set is both straightforward and has good performance. We also remove stop words, whose translations are more unstable, from the anchor set. Neither of the two modifications shows an improvement over the full anchor set. One possible explanation is that the translation qualities vary across anchors and thus we cannot predict the quality of sampled anchors.
@@ Appendix A Number of closest anchors
In addition to using all (6731) parallel anchors, we consider only the top-$k$ ($k\in\{1,10,50,100\}$) closest anchors of each word. We identify the optimal value for $k$ closest anchors based on zero-shot performance on German and Chinese portions of the Amazon Reviews Corpus (C.1). Table 3 shows results for different $k$ values.
@@ Appendix B Total number of anchors
Following Moschella et al. (2023), we randomly sample a subset of the parallel anchors ($|\boldsymbol{A}|\in\{500,3000\}$), and exclude stop words from the anchor set. Table 4 shows zero-shot performance on German and Chinese portions of the Amazon Reviews Corpus (C.1).

### v2

@@ 3 MoSECroT Task Setting
The task setting is straightforward: given a PLM of a high-resource language (regarded as the source language) and static word embeddings of another language (low-resource and regarded as the target language), the goal is to achieve zero-shot transfer by using the target language embeddings directly with the source language model via embedding layer stitching. This can be done by first applying an alignment between the source and target embedding spaces and subsequently swapping the embedding matrices of the PLM.
@@ 4 Methodology
Following Moschella et al. (2023), we build relative representations (RRs) for each token in the embedding space based on their similarities with anchor tokens in the respective language. Specifically, we compute the cosine similarity of the embedding of each token with the embedding of each anchor token. This computation is done in the embedding spaces of the source and target languages respectively. For example, in the source language, the similarity between token $x_{i}$ and anchor $a_{j}$ is calculated as follows:
@@
Note that the relative representation is sensitive to the order of the anchors, so the relative representation for each token is computed with the anchors in the same order. This computation results in a matrix $\boldsymbol{R}^{s}\in\mathbb{R}^{|V^{s}|\times|A|}$ of source language embeddings and a matrix $\boldsymbol{R}^{t}\in\mathbb{R}^{|V^{t}|\times|A|}$ of target language embeddings, where $|V^{s}|$ (resp. $|V^{t}|$) is the source-language (resp. target-language) vocabulary size and $|A|$ is the number of parallel anchors.
@@
The obtained relative representations are vectors in $\mathbb{R}^{|A|}$ for both source and target languages. This dimension does not suit the hidden dimension of the Transformer body of the source PLM. Therefore, we propose to map the relative representations of both source and target languages back to $\mathbb{R}^{D}$, which is the same as the dimension of $\boldsymbol{E}^{s}$. Given $\boldsymbol{E}^{s}$ and $\boldsymbol{R}^{s}$ for source language (resp. $\boldsymbol{E}^{t}$ and $\boldsymbol{R}^{t}$ for target language), we compute the transformed embedding of any token $x_{i}$ from the source language (resp. any token $y_{i}$ from the target language) as follows:
@@
So far we project the target-language embeddings to $\mathbb{R}^{D}$, which suits the hidden dimension of the Transformer body of the source language. We also manipulate the original token embedding matrix of the source language, where the matrix dimensions stay the same: $\boldsymbol{F}^{s}\in\mathbb{R}^{|V^{s}|\times D}$. We can simply fine-tune the model ($\boldsymbol{F}^{s}$ and the Transformer body) on the source-language train set of a downstream task and then assemble a target-language model for zero-shot transfer, without training on the target language. To do this, we only need to swap the source-language embeddings $\boldsymbol{F}^{s}$ with target-language embeddings $\boldsymbol{F}^{t}$.
@@ 5.1 Setup
Using the method proposed in §4, we obtain pairwise parallel anchors between English and each target language. The size of the anchor set varies depending on the vocabulary size of the language’s embeddings and the overlap between the English and target language lexica, which is the following for each target language: 11836 (en-de), 11395 (en-es), 7662 (en-zh), 1577 (en-fo), 2600 (en-mt), 1309 (nds), 3242 (en-sah), and 9275 (en-tt).
We evaluate the proposed method on two text classification datasets: Multilingual Amazon Reviews Corpus (Keung et al., 2020) and Taxi1500 (Ma et al., 2023). See §C for details.
Apart from the standard weighting scheme illustrated in §4, we propose two more settings: one where we apply softmax over relative representation weights (in the Embedding mapping step), and another using sparsemax (Martins and Astudillo, 2016). Compared to softmax, sparsemax produces sparse weight distributions, meaning more similarities are concentrated on fewer anchors. We conduct preliminary experiments to identify the optimal top-$k$ closest anchors $\in\{1,10,50,100\}$ and find that the results are best when using the top 50 anchors. See §A for an exploration of how different choices of $k$ influence the performance.
@@
5.2 Baselines
We compare our method against three baselines:
Logistic Regression (LR)
mBERT
We fine-tune multilingual BERT (mBERT) (Devlin et al., 2019), which is pre-trained on more than 100 languages, using the English training data, and perform zero-shot predictions directly on the target language test data.
Least squares projection (LS)
We propose a straightforward approach, inspired by embedding alignment frameworks such as VecMap (Artetxe et al., 2018), to project target language embeddings into the same space as the English PLM embeddings. Specifically, we learn a transformation matrix $\boldsymbol{W}\in\mathbb{R}^{D^{t}\times D}$ by minimizing $||\boldsymbol{A}^{t}\boldsymbol{W}-\boldsymbol{A}^{s}||_{F}^{2}$, where $\boldsymbol{A}^{t}\in\mathbb{R}^{|A|\times D^{t}}$ is the embeddings of anchors in the target language and $\boldsymbol{A}^{s}\in\mathbb{R}^{|A|\times D}$ is the embeddings of anchors from the English PLM. We then project all target language embeddings using $\boldsymbol{W}$ and replace the BERT embedding layer with the resulting matrix.
@@ 5.3 Results
We present evaluation results of RRs with the proposed settings (§5.1) and compare them with the baselines in Tables 1 and 2. Macro $F_{1}$ is used due to class imbalance in both datasets.
We notice that the naive LS baseline is almost always beaten by the proposed method under multiple RR settings on both datasets. The only exception is nds, in Table 2, where both LS and RRs perform badly. This observation is a strong indicator that RRs can better leverage the semantic similarity encoded in different types of embeddings than LS.
Not very surprisingly, zero-shot with mBERT is effective for high-resource languages in both datasets but underperforms LR with large gaps on low-resource languages in Taxi1500. There are two possible explanations for this phenomenon. First, representations in mBERT are not well-aligned across low-resource languages. This is possibly due to data sparsity, which is observed by previous work (Wu and Dredze, 2020), where mBERT archives good performance on high-resource languages but sub-optimal performance on low-resource languages. Second, Taxi1500 is a relatively easy task: a model with good alignment across languages, especially on the word level, is expected to perform well. This argument is supported by a previous work (Liu et al., 2023), where well-aligned word embeddings achieve better zero-shot crosslingual performance than mPLMs on a wide range of languages in Taxi1500.
Although none of the RR settings outperforms mBERT on high-resource languages (as mentioned earlier, mBERT has strong crosslingual transfer ability on high-resource languages), for all five low-resource languages not seen by mBERT (mt, sah, fo, nds, tt), RRs outperform mBERT consistently, with varying margins (ranging from +0.12 for sah to +0.01 for nds). This suggests that RRs can be a promising alternative when a low-resource language is not covered by an mPLM.
@@ 6 Analysis
We try to decrease the influence of potentially noisy anchor pairs by reducing the number of anchors to 3000 and 500 (the original anchor set used during the preliminary experiments contains 6731 anchors, see §4) through random sampling, following the observation by Moschella et al. (2023) that uniform selection from an anchor set is both straightforward and has good performance. We also remove stop words, whose translations are more unstable, from the anchor set. Neither of the two modifications shows an improvement over the full anchor set (see §B for the comparison). One possible explanation is that the translation qualities vary across anchors and thus we cannot predict the quality of sampled anchors.
@@ Appendix A Number of closest anchors
In addition to using all (6731) parallel anchors, we consider only the top-$k$ ($k\in\{1,10,50,100\}$) closest anchors of each word. We identify the optimal value for $k$ closest anchors based on zero-shot performance on German and Chinese portions of the Amazon Reviews Corpus (§C.1). Table 3 shows results for different $k$ values.
@@ Appendix B Total number of anchors
Following Moschella et al. (2023), we randomly sample a subset of the parallel anchors ($|\boldsymbol{A}|\in\{500,3000\}$), and exclude stop words from the anchor set. Table 4 shows zero-shot performance on German and Chinese portions of the Amazon Reviews Corpus (§C.1).